# Hire Me or Not? Examining Language Model’s Behavior with Occupation Attributes

### v1

@@ 1 Introduction
Large language models (LLMs) have become well-known to public users due to their impressive performance across multiple tasks (Tan et al., 2023; Wang et al., 2023; Lee et al., 2023) that are scalable with model size (Kaplan et al., 2020). Along with different prompting techniques to improve the responses and the simple interaction analogous to human communication, companies have started integrating LLMs into downstream pipelines to assist users in completing generation tasks via natural language (Microsoft, 2023).
However, a known issue of language models (LMs) is the human biases traced back to the large training corpus (Bender et al., 2021; Blodgett et al., 2020; Nozza et al., 2022; Smith et al., 2022; Solaiman et al., 2019; Talat et al., 2022), which can impact the fairness of the downstream tasks (Rudinger et al., 2018; Zhao et al., 2018a; Stanovsky et al., 2019; Dev et al., 2020; Liang et al., 2021; He et al., 2021). Various methods have been proposed to mitigate human biases, for example, data augmentation using counterfactuals (Zhao et al., 2019; Maudslay et al., 2019; Zmigrod et al., 2019), adjusting model parameters (Lauscher et al., 2021; Garimella et al., 2021; Kaneko & Bollegala, 2021; Guo et al., 2022), and modify decoding step to decrease the harmful generations (Schick et al., 2021). Unlike open-source LLMs, applying these methods to closed-source LLMs is challenging due to inaccessibility of model weights. Additionally, even if one can access the model weights, fine-tuning LLMs to mitigate a certain human bias may introduce new biases (Van Der Wal et al., 2022), as demonstrated by prior works in embedding debiasing methods (Bordia & Bowman, 2019; Gonen & Goldberg, 2019; Nissim et al., 2020). Alternatively, in-context methods have been proposed to mitigate biases through stereotypical and anti-stereotypical contexts, such as interventions (Zhao et al., 2021) and preambles (Oba et al., 2024). Recent work has found that LLMs still exhibit gender biases even after removing explicit signals, such as co-occurrences of “female” and “nurse”, suggesting the measured bias is not necessarily relevant to explicit gender-associated words (Belém et al., 2023).
@@
In this work, we focus on gender stereotypes related to occupation. Particularly, we investigate LLMs’ behavior with the appearance of implicit neutral occupation-relevant attribtues. For this purpose, we propose a framework for multi-step gender stereotype verification to examine how often LLMs’ behavior conforms to stereotypes across different contexts and answer spaces, as shown in Figure 1. As human biases change along with time and environment (Kozlowski et al., 2020), we leverage the latest standard occupation classification taxonomy released by O*NET (Gregory et al., 2019) as the source of implicit neutral occupation-relevant attributes.
Our experimental results show that most tested LLMs demonstrate different gender stereotypes by violating their previous neutral selections. Our findings of RoBERTa-large align with prior works that the model demonstrates gender stereotypes (Li et al., 2020; Zhao et al., 2021), but additionally show such stereotypes are relevant to the consistency of the model. The results of GPT-3.5-turbo and Llama2-70b-chat show some gender stereotypes are analogous to humans and some contradict traditional stereotypes. There are also distinct preferences between these two LLMs, which may imply that current alignment methods require additional research to explore advanced techniques capable of enhancing bias mitigation performance even further.
@@ 2 Related Work
Repetitive co-occurrences between genders and certain occupations could perpetuate and be transmitted through natural language then forming gender biases, for example, male doctors and female nurses. Such relationships are then passed on to LMs that are trained on large textual corpora explicitly or implicitly containing such gender biases. Extensive literature has shown that gender biases exist in the input representations to pre-trained language models (PLMs) (Bolukbasi et al., 2016; Caliskan et al., 2017; Zhao et al., 2017; Garg et al., 2018; Zhao et al., 2018b; May et al., 2019; Swinger et al., 2019; Zhao et al., 2019; Chaloner & Maldonado, 2019; Bordia & Bowman, 2019; Tan & Celis, 2019; Zhao et al., 2020), and downstream tasks, for example, coreference resolution (Rudinger et al., 2018; Zhao et al., 2018a; Kurita et al., 2019), machine translation (Vanmassenhove et al., 2018; Stanovsky et al., 2019; Cho et al., 2019), textual entailment (Sap et al., 2020; Dev et al., 2020), and so on (Tatman, 2017; Kiritchenko & Mohammad, 2018; Park et al., 2018; Sheng et al., 2019; Lu et al., 2020).
Some recent works have focused on probing models’ behavior via alternating the input (Wallace et al., 2019; Gardner et al., 2020; Sheng et al., 2020; Emelin et al., 2021; Ye & Ren, 2021; Schick & Schütze, 2021; Oba et al., 2024), as well as via underspecified questions (Li et al., 2020; Zhao et al., 2021).
A range of recent works investigates human biases in LLMs. Acerbi & Stubbersfield (2023) use transmission chain-like methodology to reveal that ChatGPT-3 shows biases analogous to humans for stereotypical content over other content. Gupta et al. (2024) find that LLMs are deeply biased and suggest they manifest implicit stereotypical and often erroneous presumptions when taking on a persona. Wan et al. (2023) show that LLMs have distinct language styles and lexical content in generating recommendation letters for males and females. Belém et al. (2023) demonstrate that measured gender bias is not necessarily due to the explicit signals, suggesting the implicit factors that contribute to the biased behavior of LLMs. Kotek et al. (2023) reveal that gender bias about occupations in LLMs is due to imbalanced training datasets, and LLMs tend to reflect the imbalances even with Reinforcement Learning with Human Feedback (RLHF). Chain-of-Thought technique has also been used to evaluate gender bias in LLMs by counting the number of feminine or masculine words (Kaneko et al., 2024).
Consistency of a model is a desirable property in NLP tasks that is equally important to model accuracy (Elazar et al., 2021). There are many prior works exploring consistency of PLMs for question answering (Rajpurkar et al., 2016; Ribeiro et al., 2019; Alberti et al., 2019; Asai & Hajishirzi, 2020; Kassner et al., 2021), robust evaluation(Li et al., 2019), natural language inference (Camburu et al., 2018; 2020), and more (Du et al., 2019).
@@ 3 Multi-step Gender Stereotypes Verification
In order to provide background information conducive to multi-step question answering, we integrated structured human knowledge about occupations from authoritative labor statistics. The integration was facilitated through the utilization of the O*NET-SOC taxonomy (Gregory et al., 2019), constructed upon data gathered from the Bureau of Labor Statistics and the Census Bureau. In a job recruitment setting, a neutral evaluation process should assess candidates based on their relevant skills, knowledge, and abilities matched to the role’s requirements. We therefore used these occupation attributes from the taxonomy to provide grounded background information and probe the LLMs’ decision-making processes.
@@ 3.1 Dataset Construction
Our dataset focuses on two subject categories: male and female, aiming to investigate potential gender biases regarding various occupation titles. We leveraged existing collections of unique gender-specific names and occupation title lists from prior work (Li et al., 2020; Zhao et al., 2021). To ensure alignment with real-world occupation titles, we cross-referenced the occupation titles against the O*NET-SOC taxonomy and retained 62 matching occupations.
@@
Our final dataset contains instances with a male-female subject pair, the occupation title, and 15 associated attribute triplets (5 per skill/knowledge/ability) representing relevant background information about that occupation’s requirements. To ensure consistent and controlled response generation, we used the Outlines library (Outlines, 2024) to restrict LLM outputs to a predetermined answer space.
@@ 3.2 Extracting Background Attribute Associations
To begin, we probed the relevant attributes of each occupation (skills, knowledge, abilities) associated with each subject in a gender pair using LLMs. Following the format of the UnQover dataset (Li et al., 2020), each instance contained a base context paragraph $p$ and an underspecified question $q$:
@@
In Step 1, as illustrated in Figure 2, each LLM outputs Yes/No to indicate whether the subject possesses a particular attribute. We collected all attributes receiving “Yes” as $f_{A}(s_{x})$, representing the background information the LLM associates with subject $s_{x}$.
@@ 4 Results
  • RoBERTa-large (Liu et al., 2019) fine-tuned on SQuAD v2.0 (Rajpurkar et al., 2018)
  • GPT-3.5-turbo (OpenAI, 2021)
  • Llama2-70b-chat (Touvron et al., 2023)
@@ 5.1 Limitation
Our work is limited to investigating gender biases and stereotypes in English, a morphologically limited language. Recent studies have found gender biases existing in LLMs for different languages (Malik et al., 2022; Névéol et al., 2022; Kaneko et al., 2022; Anantaprayoon et al., 2023; Levy et al., 2023). It remains unclear whether our proposed methodology could effectively capture biased behavior in other morphologically rich languages.
Moreover, we focused solely on gender biases in this work. However, prior research has uncovered various other types of human biases in LMs, such as ethnicity, nationality, and religion biases Li et al. (2020); Zhao et al. (2021). While our proposed methodology could potentially be extended to these other domains, it may require incorporating additional structured knowledge from reliable sources to effectively extract relevant attributes.
Furthermore, we only considered gender-specific names in our work. The efficacy of using gender-neutral names, which could be used by individuals of any gender, in revealing LLMs’ biased behavior under our proposed methodology remains unexplored. Additionally, our work only addresses binary gender biases, whereas non-binary gender biases have also been explored in recent literature (Cao & Daumé III, 2020; Dev et al., 2021).

### v2

# Damin Zhang, Yi Zhang, Geetanjali Bihani, Julia Rayz

@@ 1 Introduction
Large language models (LLMs) have become well-known to public users due to their impressive performance across multiple tasks [1, 2, 3] that are scalable with model size [4]. Along with different prompting techniques to improve the responses and the simple interaction analogous to human communication, companies have started integrating LLMs into downstream pipelines to assist users in completing generation tasks via natural language [5].
However, a known issue of language models (LMs) is the human biases traced back to the large training corpus [6, 7, 8, 9, 10, 11], which can impact the fairness of the downstream tasks [12, 13, 14, 15, 16, 17]. Various methods have been proposed to mitigate human biases, for example, data augmentation using counterfactuals [18, 19, 20], adjusting model parameters [21, 22, 23, 24], and modify decoding step to decrease the harmful generations [25]. Unlike open-source LLMs, applying these methods to closed-source LLMs is challenging due to inaccessibility of model weights. Additionally, even if one can access the model weights, fine-tuning LLMs to mitigate a certain human bias may introduce new biases [26], as demonstrated by prior works in embedding debiasing methods [27, 28, 29]. Alternatively, in-context methods have been proposed to mitigate biases through stereotypical and anti-stereotypical contexts, such as interventions [30] and preambles [31]. Recent work has found that LLMs still exhibit gender biases even after removing explicit signals, such as co-occurrences of “female" and “nurse", suggesting the measured bias is not necessarily relevant to explicit gender-associated words [32].
@@
In this work, we focus on gender stereotypes related to occupation. Particularly, we investigate LLMs’ behavior with the appearance of implicit neutral occupation-relevant attribtues. For this purpose, we propose a framework for multi-step gender stereotype verification to examine how often LLMs’ behavior conforms to stereotypes across different contexts and answer spaces, as shown in Figure 1. As human biases change along with time and environment [33], we leverage the latest standard occupation classification taxonomy released by O*NET [34] as the source of implicit neutral occupation-relevant attributes.
Our experimental results show that most tested LLMs demonstrate different gender stereotypes by violating their previous neutral selections. Our findings of RoBERTa-large align with prior works that the model demonstrates gender stereotypes [35, 30], but additionally show such stereotypes are relevant to the consistency of the model. The results of GPT-3.5-turbo and Llama2-70b-chat show some gender stereotypes are analogous to humans and some contradict traditional stereotypes. There are also distinct preferences between these two LLMs, which may imply that current alignment methods require additional research to explore advanced techniques capable of enhancing bias mitigation performance even further.
@@ 2 Related Work
Repetitive co-occurrences between genders and certain occupations could perpetuate and be transmitted through natural language then forming gender biases, for example, male doctors and female nurses. Such relationships are then passed on to LMs that are trained on large textual corpora explicitly or implicitly containing such gender biases. Extensive literature has shown that gender biases exist in the input representations to pre-trained language models (PLMs) [36, 37, 38, 39, 40, 41, 42, 18, 43, 27, 44, 45], and downstream tasks, for example, coreference resolution [12, 13, 46], machine translation [47, 14, 48], textual entailment [49, 15], and so on [50, 51, 52, 53, 54].
Some recent works have focused on probing models’ behavior via alternating the input [55, 56, 57, 58, 59, 60, 31], as well as via underspecified questions [35, 30].
A range of recent works investigates human biases in LLMs. [61] use transmission chain-like methodology to reveal that ChatGPT-3 shows biases analogous to humans for stereotypical content over other content. [62] find that LLMs are deeply biased and suggest they manifest implicit stereotypical and often erroneous presumptions when taking on a persona. [63] show that LLMs have distinct language styles and lexical content in generating recommendation letters for males and females. [32] demonstrate that measured gender bias is not necessarily due to the explicit signals, suggesting the implicit factors that contribute to the biased behavior of LLMs. [64] reveal that gender bias about occupations in LLMs is due to imbalanced training datasets, and LLMs tend to reflect the imbalances even with Reinforcement Learning with Human Feedback (RLHF). Chain-of-Thought technique has also been used to evaluate gender bias in LLMs by counting the number of feminine or masculine words [65].
Consistency of a model is a desirable property in NLP tasks that is equally important to model accuracy [66]. There are many prior works exploring consistency of PLMs for question answering [67, 68, 69, 70, 71], robust evaluation[72], natural language inference [73, 74], and more [75].
@@ 3 Multi-step Gender Stereotypes Verification
In order to provide background information conducive to multi-step question answering, we integrated structured human knowledge about occupations from authoritative labor statistics. The integration was facilitated through the utilization of the O*NET-SOC taxonomy [34], constructed upon data gathered from the Bureau of Labor Statistics and the Census Bureau. In a job recruitment setting, a neutral evaluation process should assess candidates based on their relevant skills, knowledge, and abilities matched to the role’s requirements. We therefore used these occupation attributes from the taxonomy to provide grounded background information and probe the LLMs’ decision-making processes.
@@ 3.1 Dataset Construction
Our dataset focuses on two subject categories: male and female, aiming to investigate potential gender biases regarding various occupation titles. We leveraged existing collections of unique gender-specific names and occupation title lists from prior work [35, 30]. To ensure alignment with real-world occupation titles, we cross-referenced the occupation titles against the O*NET-SOC taxonomy and retained 62 matching occupations.
@@
Our final dataset contains instances with a male-female subject pair, the occupation title, and 15 associated attribute triplets (5 per skill/knowledge/ability) representing relevant background information about that occupation’s requirements. To ensure consistent and controlled response generation, we used the Outlines library [76] to restrict LLM outputs to a predetermined answer space.
@@ 3.2 Extracting Background Attribute Associations
To begin, we probed the relevant attributes of each occupation (skills, knowledge, abilities) associated with each subject in a gender pair using LLMs. Following the format of the UnQover dataset [35], each instance contained a base context paragraph $p$ and an underspecified question $q$:
@@
In Step 1, as illustrated in Figure 2, each LLM outputs Yes/No to indicate whether the subject possesses a particular attribute. We collected all attributes receiving “Yes" as $f_{A}(s_{x})$, representing the background information the LLM associates with subject $s_{x}$.
@@ 4 Results
  • RoBERTa-large [77] fine-tuned on SQuAD v2.0 [78]
  • GPT-3.5-turbo [79]
  • Llama2-70b-chat [80]
@@ 5.1 Limitation
Our work is limited to investigating gender biases and stereotypes in English, a morphologically limited language. Recent studies have found gender biases existing in LLMs for different languages [81, 82, 83, 84, 85]. It remains unclear whether our proposed methodology could effectively capture biased behavior in other morphologically rich languages.
Moreover, we focused solely on gender biases in this work. However, prior research has uncovered various other types of human biases in LMs, such as ethnicity, nationality, and religion biases [35, 30]. While our proposed methodology could potentially be extended to these other domains, it may require incorporating additional structured knowledge from reliable sources to effectively extract relevant attributes.
Furthermore, we only considered gender-specific names in our work. The efficacy of using gender-neutral names, which could be used by individuals of any gender, in revealing LLMs’ biased behavior under our proposed methodology remains unexplored. Additionally, our work only addresses binary gender biases, whereas non-binary gender biases have also been explored in recent literature [86, 87].